OpenStudy (calculusxy):

Can someone help me please make a COMPLEX word problem for my project? I will surely give a medal. Info will be provided.

OpenStudy (calculusxy):

@e.mccormick

OpenStudy (anonymous):

Make a related rates problem :P

OpenStudy (anonymous):

Or a volumes problem

OpenStudy (calculusxy):

 3 cups of peeled  3 cups of diced pumpkin  4 3/4 Bunch of fresh spinach (chopped)  1/2 lb. baby shrimp (peeled and steamed)  4 tablespoons vegetable oil  3/4 cup of red chili  1 1/4 tablespoons of sugar  1 4/5 teaspoons of salt  1/2 teaspoon coriander powder  3/4 teaspoon of cumin powder  2/3 teaspoon of turmeric powder  1/4 teaspoons of chili powder  2 tablespoons of grounded/grated coconut

OpenStudy (calculusxy):

@TuringTest

OpenStudy (calculusxy):

Please help me! I need this!

OpenStudy (turingtest):

what's with the recipes?

OpenStudy (anonymous):

Now I'm confused lol, what are the ingredients for o_O, are we cooking something?

OpenStudy (calculusxy):

Yes @TuringTest @iambatman

OpenStudy (turingtest):

is this a language barrier we are having @calculusxy ? how is your English?

OpenStudy (calculusxy):

I meant on the first bullet 3 cups of peeled and diced pumpkin

OpenStudy (turingtest):

what about it?

OpenStudy (calculusxy):

I need to produce a word problem (which is complex) and related to the word problem. But I have no idea where to start.

OpenStudy (turingtest):

and you want it to be related to diced pumpkins?

OpenStudy (calculusxy):

No. The whole recipe.

OpenStudy (turingtest):

that seems tedious, pointless, and intractable. I suggest you try to think of a more elegant kind of problem with much fewer factors.

OpenStudy (calculusxy):

But this is what I am stuck with. Please help me.

OpenStudy (calculusxy):

On the prompt it says "Write one word problem from your recipe".

OpenStudy (anonymous):

Just make something up, shouldn't be too hard, I have no interest in helping with this, it's too time consuming.

OpenStudy (turingtest):

ooook so we don't have to use the all the ingredients, per se

OpenStudy (calculusxy):

Not really.

OpenStudy (turingtest):

this is calc 1 ?

OpenStudy (calculusxy):

No. Just a middle school-er word problem.

OpenStudy (anonymous):

OoOoOoOoOo

OpenStudy (turingtest):

lol your username confused both of us XD ok we can do something here

OpenStudy (anonymous):

^

OpenStudy (calculusxy):

Sorry about that :/ I like Calc, that's why.

OpenStudy (turingtest):

you can do things like "if suzie only has only 2.5 cups of peeled and diced pumpkins, what must the ratios for the other ingredients be to keep the proportions the same?"

OpenStudy (turingtest):

that should keep a middle-schooler busy... though it's not an assignment I'd want to do at that age lol

OpenStudy (calculusxy):

Well I wasn't really thinking about that complex. I hope you understand what i mean. I was thinking if I can make a problem with Part A and Part B, where Part B depends on the answer from Part A. If you can help, I will truly appreciate it :)

OpenStudy (anonymous):

So a substitution problem?

OpenStudy (anonymous):

Take ingrediant a put in b etc.

OpenStudy (anonymous):

Ingredient*

OpenStudy (calculusxy):

No like where you have an answer from the question from Part A in which the answer to Part B will be depended upon. So if you get it wrong, then Part B will be wrong. Vice-versa.

OpenStudy (kc_kennylau):

So you're supposed to create a problem based on the recipe?

OpenStudy (turingtest):

a) Suzie has only 2 cups of diced pumpkins when she needs 3. How many tablespoons of sugar must she use to keep the same proportions? b) after making the cake (or whatever it is), 4 more people come. 2 want 1/3 of a cake, and and the other two each want a whole. How many more tablespoons of sugar does she need?

OpenStudy (turingtest):

something like that? I don't know about "vice versa", as I can't see how getting b) wrong could affect a)

OpenStudy (calculusxy):

"Vice-versa" means "the other way around".

OpenStudy (turingtest):

yeah, but you said "if they get a) wrong they get b) wrong, and vice versa" but how can the first answer depend on the second *and* the second depend on the first?

OpenStudy (turingtest):

I can only see a) being necessary to solve b)

OpenStudy (calculusxy):

I meant that if the answer to part a is wrong then the answer to part b will be wrong. but vice-versa meant that if the answer to part a is correct then the answer to part b will be correct also.

OpenStudy (turingtest):

I see, so does mine fit the bill?

OpenStudy (calculusxy):

Yeah ( I would say) because I didn't do the problem yet. can you please help me with solving the problem? i will truly appreciate your help towards my project.

OpenStudy (turingtest):

actually I should modify it a bit so that b) depends entirely on a), so let's make it a) Suzie has only 2 cups of diced pumpkins when she needs 3. How many tablespoons of sugar must she use to keep the same proportions? b) the smaller cake she makes with those proportions become popular, so 4 more people come wanting cakes made with the same proportion ofingredients. 2 want 1/3 of a cake, and and the other two each want a whole. How many more tablespoons of sugar does she need?

OpenStudy (turingtest):

part a) is a rather direct proportion problem that can be solved by looking at the ratios 2/3=x/(1+1/4) solve for x

OpenStudy (turingtest):

part b) can be solved by multiplying the total number of desired pies by x

OpenStudy (calculusxy):

How did you get 1/4?

OpenStudy (turingtest):

recipe calls for 3 cups of diced pumpkin 1 1/4 tablespoons sugar

OpenStudy (turingtest):

so if we only have 2 cups dice pumpkin pumpkin available/pumpkin called for=sugar available/sugar called for 2/3=x/(1 1/4)

OpenStudy (calculusxy):

\[\frac{ 2 }{ 3 }=\frac{ x}{ 1.25 }\] Is this how it is to be set up?

OpenStudy (turingtest):

yes

OpenStudy (calculusxy):

If I do solve for x, it comes up as to a huge quotient. 0.208333333....

OpenStudy (turingtest):

no, you did something wrong

OpenStudy (calculusxy):

Okay so I first have to multiply 1.25*2 right?

OpenStudy (turingtest):

where'd the 3 go?

OpenStudy (calculusxy):

Well I thought that we were solving for x.

OpenStudy (turingtest):

we are

OpenStudy (calculusxy):

So isn't it supposed to be: \[1.25 \times 2 = 3x\]

OpenStudy (turingtest):

you could do it that way, though it's rather pointless to move the 3 over; you have to move it back where it was again anyway

OpenStudy (calculusxy):

Can you show what you mean by that @TuringTest

OpenStudy (anonymous):

1.25 * 2/3 = x, that's what he means.

OpenStudy (calculusxy):

\[\frac{ 1.25 \times 2 }{ 3 } = \frac{ 2.5 }{ 3 }=0.8333...\]

OpenStudy (turingtest):

that sounds reasonable, I'm not going to break out the calculator... since i don't own one

OpenStudy (calculusxy):

So how should I put that as my answer?

OpenStudy (turingtest):

i'd keep it as a fraction, but that's just me

OpenStudy (anonymous):

You can say approximate answer, so it can be 5/6.

OpenStudy (calculusxy):

Okay. Now up to the second problem.

OpenStudy (turingtest):

let me simplify the second part just a bit to avoid confusion, though it makes it a bit easier b) the smaller cake she makes with those proportions become popular, so 4 more people order cakes made with the same proportion of ingredients. 2 want 1/2 of a cake, and and the other two each want a whole. How many more tablespoons of sugar does she need?

OpenStudy (turingtest):

so how many pies does she need to make?

OpenStudy (calculusxy):

so you want me to use 2.5/3 for this problem also?

OpenStudy (turingtest):

why would I want that?

OpenStudy (turingtest):

oh you mean for the sugar, yeah

OpenStudy (calculusxy):

okay so from what i can understand, people want altogether 2 cakes right?

OpenStudy (turingtest):

no, 4 people order. two each want a half a cake, and two each want a whole

OpenStudy (calculusxy):

1/2 + 1/2 = 1 1 + 1= 2 sorry all together 3 cakes right?

OpenStudy (turingtest):

yes

OpenStudy (calculusxy):

proportions right?

OpenStudy (turingtest):

not really... we already have the sugar required for each cake is 5/6 tablespoons so if I make 3 cakes how many tablespoons do I need?

OpenStudy (calculusxy):

5/6*3?

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

@TuringTest Good job! :P

OpenStudy (turingtest):

Thanks @iambatman :D

OpenStudy (calculusxy):

Okay one more question, how did you get 5/6 (like multiply both by 2)?

OpenStudy (turingtest):

5/6 is 2.5/3 because... 2.5=5/2 so 2.5/3=(5/2)/3=5/6

OpenStudy (calculusxy):

Thanks a lot!!!! @TuringTest

OpenStudy (turingtest):

happy to help!

OpenStudy (kainui):

Yeah but where's the square root of negative 1?!

OpenStudy (turingtest):

um.... not in this problem thank fully. I don't like imaginary cakes

OpenStudy (calculusxy):

It wasn't really a cake. It was an Asian food called: "Saag Kumro"

OpenStudy (turingtest):

Oh nice. If I ever feel like making it I'll refer to this post :) See ya 'round!

OpenStudy (calculusxy):

You too. TY

OpenStudy (turingtest):

I guess I should have figured with the shrimp and chili lol XD